OpenStudy (anonymous):

Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (whpalmer4):

You can decompose to x and y components by \[x = r \cos \theta\]\[y = r \sin\theta\]

OpenStudy (anonymous):

Where r is what? the 200/100 and theta is 60 and 170?

OpenStudy (whpalmer4):

R is magnitude of vector, so 200 or 100 (you do this procedure twice)

OpenStudy (whpalmer4):

Theta is the angle of course

OpenStudy (anonymous):

ok, so x=200cos60 and y=100cos170 and what would this be finding? the terminal point of the vector?

OpenStudy (whpalmer4):

Then sum x components to get x component of new vector, and sum y components to get y component of new vector

OpenStudy (anonymous):

you can easily see that the angle between the forces is 170-60 =110 degrees. You have the magnitude of the two forces and the angle between them.. now use this formula to get the resultant,  and for direction, find the angle the Resultant R makes with the first force using,  Then you can say that the resultant makes an angle